2 point charges q1=+2microC at (0,0) and q2=-3microC
at (100,0) At wht pt. on x axis will the absolute potential =0?
ANs :- x=40,-200

Grade:12

1 Answers

Vikas TU
6 years ago
Dear Student,
Let at x distance from origin the net Potential is zero.
Then the resultant potential is =>
K*2/x + K*-3/(100-x) = 0
2/x = 3/(100-x)
Solving we get,
3x = 200 – 2x
5x = 200
x = 40 units.
Hence from 40 units distnce from the origin the net potential would be zero.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)